OpenStudy (anonymous):

I say x^2/9 +y^2/100 = 100 is an ellipse. Do you agree? Choice are ellipse, hyperbole, parabola, circle

OpenStudy (anonymous):

yes it is an ellipse because of the + sign

OpenStudy (anonymous):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]